### v1

# Limits of Absoluteness of Observed Events in Timelike Scenarios: A No-Go Theorem

Abstract
Wigner’s Friend–type paradoxes challenge the assumption that events are absolute — that when we measure a system, we obtain a single result, which is not relative to anything or anyone else. These paradoxes highlight the tension between quantum theory and our intuitions about reality being observer-independent. Building on a recent result that developed these paradoxes into a no-go theorem, namely the Local Friendliness Theorem, we introduce the Causal Friendliness Paradox, a time-ordered analogue of it. In this framework, we replace the usual locality assumption with Axiological Time Symmetry (ATS), and show that, when combined with the assumptions of Absoluteness of Observed Events (AOE), No Retrocausality (NRC), and Screening via Pseudo Events (SPE), we obtain a causal inequality. We then show that quantum mechanics violates this inequality and is therefore incompatible with at least one of these assumptions. To probe which assumption might be incompatible, we then examine whether AOE in its entirety is essential for this no-go result. We propose a weaker, operational form of AOE that still leads to inequalities which quantum mechanics violates. This result shows that even under relaxed assumptions, quantum theory resists reconciliation with classical notions of absolute events, reinforcing the foundational significance of Wigner’s Friend-type paradoxes in timelike scenarios.
I Introduction
One of the most unsettling quantum paradoxes proposed so far is that of Wigner’s Friend, which asks the question of what happens when, as should be valid in standard quantum mechanics, we treat human observers as quantum systems [1]. While originally simply troubling to our intuitions, in the same way as Schrödinger’s Cat [2] or the EPR paradox [3], recent work has shown that, by combining an extension of the Wigner’s Friend scenario with certain intuitive assumptions (that any event an observer sees is real and absolute, and so alternative, or “parallel”, versions of such observations do not occur simultaneously; that our choices are free and not influenced by other factors; and that no information can travel faster than $c$, the speed of light in a vacuum), we obtain predictions for the values of certain observables in the scenario which differ from the values predicted by standard quantum mechanics [4]. While these scenarios are practically impossible to perform in the real world, given they involve being able to unitarily-undo the process of a human observer measuring a quantum system, it is deeply interesting that we are able to obtain such a contradiction, especially given the assumptions are arguably weaker than those used in similar no-go results, such as Bell’s Theorem [5, 6, 7, 8, 9, 10, 11, 12, 13] or the Kochen-Specker Theorem [14, 15].
However, it is still an open question whether any of the assumptions behind the so-called Local Friendliness inequalities [4] in this Extended Wigner’s Friend scenario can be weakened any further. In this paper, we show that the assumptions can be weakened, in two ways.
Firstly, we show that we can replace the Locality assumption (the assumption that no information travels faster than $c$) in this standard Extended Wigner’s Friend scenario with an assumption of time-symmetry in a time-ordered version of the scenario - i.e., by making the scenario causally rather than spatially separated, we can weaken the assumptions by just assuming the choices in the first part of the experiment always affect the probabilities of the second part of the experiment in the same way, regardless of which part of the experiment is done first and which is done second.
Secondly, we show that in this time-ordered scenario, we can also weaken the assumption of Absoluteness of Observed Events (that any event an observer sees is real and absolute, meaning no “parallel” versions of such an observation exist with different measurement results), by an operational analogy of the weakening of the ontological assumption in Price’s no-go result [16] by Leifer and Pusey [17].
One counterargument against the intuitiveness and viability of the assumptions used to generate the Local Friendliness inequalities is that Absoluteness of Observed Events, and Unitary Undoing of an Observer’s Measurement, are fundamentally incompatible, regardless of any other assumptions. Showing the inequality still holds for the weakened form of AOE implies that this counterargument fails to account for the entirety of the peculiarity of quantum mechanics’ ability to violate these Local (or here, Causal) Friendliness inequalities.
This paper is laid out as follows. In Section II, we first go through the original (Section II.1) and extended (Section II.2) Wigner’s Friend thought experiments, with special focus for the Extended Wigner’s Friend thought experiment on the assumptions required to achieve a contradiction with standard quantum mechanics.
In Section III, we extend these results to a new paradox: the Causal Friendliness Paradox. In Section III.1, we describe the (hypothetical) experimental scenario for the Causal Friendliness Paradox to occur, before in Section III.2 giving the assumptions required for such a scenario to give predictions contradicting standard quantum mechanics (a contradiction we formally prove in Subsection III.3).
In Section IV, we expand on our result from Section III, by seeing to what extent we can weaken the Absoluteness of Observed Events assumption common to both our Causal Friendliness Paradox and the Extended Wigner’s Friend Paradoxes. We do this by looking at a result by Leifer and Pusey [17], where they weakened the $\psi$-ontic reproducibility assumption in Price’s 2007 no-go result on time-symmetry and retrocausality [16], to instead just require $\lambda$-mediation. In Section IV.1, we review how Leifer and Pusey weakened the $\psi$-ontic reproducibility assumption into $\lambda$-mediation. In Section IV.2, we then analogise this weakening at an operational, rather than ontological, level, to enable us to weaken the Absoluteness of Observed Events assumption while still being able to generate an inequality that quantum mechanics violates. However, in Section IV.3, we then show that we cannot weaken this assumption any further (i.e., specifically remove the Absoluteness of the Pseudo Events - the Friends measuring their system - which may or may not be unitarily “rewound” by the Wigners) without losing the ability to generate such a contradiction, with either quantum correlations or even Popescu and Rohrlich’s Boxworld’s correlations [18].
Finally, in Section V, we discuss our results: in Section V.1, we compare our results to previous work; and in Section V.2, we discuss the implications of our work for the study of Wigner’s Friend-style paradoxes, and quantum foundations more broadly.
II Wigner’s Friend Scenarios and their implications
II.1 The Wigner’s Friend paradox
One of the peculiarities implied by the measurement problem [19, 20, 21] was illustrated by Wigner in his eponymous thought experiment, the Wigner’s Friend paradox [1]. The paradox involves a scenario with two observers, Wigner and his Friend, each capable of performing specific operations on the quantum systems assigned to them. While the Friend remains inside a laboratory $\mathbf{L}$, isolated from the external environment and measuring a quantum system $\mathbf{S}$ contained within, Wigner stays outside and treats the entire lab as a single quantum system.
The quantum system $\mathbf{S}$ inside the lab is described by a wave function
$\left|\psi\right\rangle_{S}=\frac{\left|\uparrow\right\rangle_{S}+\left|\downarrow\right\rangle_{S}}{\sqrt{2}},$ (1)
where $\{\left|\uparrow\right\rangle_{S},\left|\downarrow\right\rangle_{S}\}$ are eigenstates of spin of a system in some arbitrary direction. Therefore, if the Friend measures the system, she gets an outcome corresponding to either $\left|\uparrow\right\rangle_{S}$ or $\left|\downarrow\right\rangle_{S}$. However, from Wigner’s perspective, the entire laboratory behaves as a large quantum system. This includes not only the system being measured by the Friend, but also the measurement device, whose corresponding quantum state can be either $\left|\uparrow\right\rangle_{D}$ or $\left|\downarrow\right\rangle_{D}$ depending on the state of the system, as well as the Friend, whose quantum state can similarly be $\left|\uparrow\right\rangle_{F}$ or $\left|\uparrow\right\rangle_{F}$. Altogether, there exists a linear map $f$ from the wave function of the measured system to the combined wave function of the full laboratory, given by:
$f:\left\{\begin{aligned} &\left|\uparrow\right\rangle_{S}\rightarrow\left|\uparrow\right\rangle_{S}\otimes\left|\uparrow\right\rangle_{D}\otimes\left|\uparrow\right\rangle_{F}\equiv\left|\uparrow\right\rangle_{L}\\ &\left|\downarrow\right\rangle_{S}\rightarrow\left|\downarrow\right\rangle_{S}\otimes\left|\downarrow\right\rangle_{D}\otimes\left|\downarrow\right\rangle_{F}\equiv\left|\downarrow\right\rangle_{L}\end{aligned}\right$. (2)
Therefore, if the initial state of the system is $\left|\psi\right\rangle_{S}$, then after the Friend’s measurement, the wave function of the lab (from Wigner’s perspective) takes the form,
$\left|\psi\right\rangle_{W}=\frac{\left|\uparrow\right\rangle_{L}+\left|\downarrow\right\rangle_{L}}{\sqrt{2}}$ (3)
As the lab is isolated from the environment, Wigner has no way to know the result of the measurement that the Friend has already performed. Therefore, to Wigner, the state of the system remains as a probabilistic mixture of $\left|\uparrow\right\rangle_{S}$ and $\left|\downarrow\right\rangle_{S}$, which he gets by tracing out the device and Friend’s wave function. In contrast, the Friend has already obtained a definite outcome (so to them, the state of the system is either in the pure state $\left|\uparrow\right\rangle_{S}$ or the pure state $\left|\downarrow\right\rangle_{S}$). This apparent contradiction between Wigner’s and his Friend’s perceptions (and descriptions) of a single event highlights the incommensurability of quantum theory in accounting for observer-independent facts.
However, on opening the Friend’s lab, Wigner (and the wider environment) also becomes entangled with the lab, and so the Friend’s system. This means, on opening the lab, the state of the system (from Wigner’s perspective) also becomes a single eigenstate (either the pure state $\left|\uparrow\right\rangle_{S}$, or the pure state $\left|\downarrow\right\rangle_{S}$). Therefore, despite the difference in the state that Wigner and the Friend each associate to the Friend’s measured system while the lab is closed, there is no point at which these two different states give predictions which could observably contradict. One could simply dismiss any apparent contradiction by asserting that unperformed experiments have no results [22]. This led to the question of what assumptions would be necessary for us to adapt the scenario in such a way that experimentally-testable predictions derived from these assumptions contradicted the predictions of quantum theory.
II.2 Extended Wigner Friend Scenarios
Figure 1: The extended Wigner’s Friend Scenario for the Local Friendliness no-go theorem.
Recently, a flurry of works has emerged [23, 24, 25, 26, 27, 28, 29], building upon the original Wigner’s Friend paradox [1] and its subsequent generalisations [30], and exploring various foundational issues related to these scenarios. Frauchiger and Renner [25] first proposed an extension of the Wigner’s Friend paradox in which the quantum probability assignments are shown to be logically inconsistent with the assumption of observer-independent facts. However, it was only in Brukner’s later version [23] that such inconsistencies were used to derive differing, testable predictions. Although these works received significant attention, their conclusions rely on counterfactual reasoning—an aspect which arguably weakened their claims. Bong et al later put forward a stronger no-go theorem [4], which, rather than challenging the notion of observer-independent facts involving counterfactual scenarios, directly questions the absoluteness of already observed events.
Inspired by Brukner’s proposal [23], Bong et al  [4] extended the Wigner’s Friend thought experiment to the scenario where two observers (Charlie and Debbie) each possess one part of a non-trivially correlated pair (i.e., in quantum theory, an entangled state). These two observers are in labs $\mathbf{L}_{C}$ and $\mathbf{L}_{D}$ respectively. Further, there exist two more observers, Alice and Bob, who can perform any arbitrary operation on lab $\mathbf{L}_{C}$ and $\mathbf{L}_{D}$ respectively. As they each have this power of performing measurements over one of the labs (and so one of the observers), they are referred to as super-observers. This is pictorially depicted in Fig. 1.
In this scenario, Charlie measures his system (staying inside his lab), obtaining output $c\in\{0,1\}$. This measurement process is describable by a unitary operation $\mathcal{U}_{C}$ on the lab $\mathbf{L}_{C}$. After Charlie’s measurement, Alice chooses $x\in\{0,1,2,..\}$ (her input) at random, acting depending on the value of $x$. If $x=0$, then Alice opens lab $\mathbf{L}_{C}$ and asks Charlie about his measurement result, setting her output $a\in\{0,1\}$ equal to Charlie’s output $c$. In other cases (i.e., if $x=1,2,..$), Alice applies $\mathcal{U}_{C}^{\dagger}$ to $\mathbf{L}_{C}$, undoing Charlie’s measurement of the system, before opening lab $\mathbf{L}_{C}$, and measuring Charlie’s system herself, setting her output $a$ as the result of this measurement. Debbie and Bob proceed in the same manner as Charlie and Alice on the other side of the experimental set-up, except Bob uses random input $y\in\{0,1,2,$...\} instead of $x$, and Debbie’s and Bob’s outcomes are denoted by $d\in\{0,1\}$ and $b\in\{0,1\}$, respectively (rather than $c$ and $a$).
Bong et al derived what they call Local Friendliness (LF) inequalities by considering the empirical probabilities $p(a,b|x,y)$ we get for this scenario when we impose the following set of assumptions:
  • Assumption 1. Absoluteness of Observed Events (AOE): An observed event is a real single event, and not relative to anything or anyone.
    This implies that there exists a joint probability distribution $p(a,b,c,d|x,y)$ such that all the empirical probabilities can be recovered from it, as
    $p(a,b|x,y)=\sum_{c,d}p(a,b,c,d|xy)\ \ \forall a,b,x,y$ (4)
  • Assumption 2. No-Superdeterminism (NSD): Any set of unitarily-erasable events on a space-like hypersurface is uncorrelated with any set of freely chosen actions subsequent to that space-like hypersurface.
    Here, this implies that the joint probability of outcomes $c$ and $d$ is independent of $x$ and $y$, i.e., $p(c,d|xy)=p(c,d)$. In the Bell scenario, this assumption is known as Statistical Independence [13] (often referred to erroneously as the “Free Will” assumption), which says that the hidden variables are conditionally independent of the choice of measurements. However, in contrast to the Bell scenario, here the hidden variable-like values $c$ and $d$ are outcomes of (unitarily-erasable) observed events.
  • Assumption 3. Locality (L): The probability of an observable event $e$ is unchanged by conditioning on a space-like-separated free choice $z$, even if it is already conditioned on other events (so long as these events are not in the future light-cone of $z$).
    For the current set-up, the locality assumption reads:
    $p(a|c,d,x,y)=p(a|c,d,x)\ \ \forall c,d,x,y$
    $p(b|c,d,x,y)=p(b|c,d,y)\ \ \forall c,d,x,y$ (5)
This assumption is similar to the Parameter Independence assumptions used in the conventional Bell scenario [31, 32] - the only difference being that here the (potentially-observed) events $c$ and $d$ replace the hidden variable $\lambda$ from the Bell scenario.
Bong et al showed that quantum theory violates the derived inequality, and from this violation they claimed: as Assumptions 2 and 3 are (supposedly) satisfied by quantum statistics, then Assumption 1 must be violated. They therefore argue that quantum theory is incompatible with the assumption of Absoluteness of Observed Events.
It is important to note that, even though the events $c$ and $d$ are treated as observed events in Bong et al ’s treatment, this is not necessarily the case: when $x\neq 0$ ($y\neq 0$), event $c$ ($d$) is unitarily erased by Alice (Bob), as if a value was never assigned to the variable. Thus, we argue that events $a$,$b$ and events $c$,$d$ cannot be treated in the same way. When $x\neq 0$ ($y\neq 0$) we therefore call $c$ ($d$) Pseudo Events, while $a$ and $b$ we refer to as Truly-Observed Events.
Figure 2: Time-ordered measurement scenario for Causal Friendliness no-go theorem.
III Causal Friendliness Paradox
III.1 The Causal Friendliness Scenario
Here, we introduce our new scenario, illustrated in Fig. 2. As in Ref. [4], Charlie and Debbie begin the experiment sealed in their respective labs. However, here they do not share an entangled state. Instead, Charlie measures his system first, with his measurement outcome denoted by $c\in\{0,1\}$. Given his lab is completely closed to the environment, we again assume this measurement process can be represented by the unitary action $\mathcal{U}_{C}$ from the point of view of the Alice, who resides outside the lab and acts as an super-observer. After Charlie’s measurement, Alice acts based on her random input $x\in\{0,1\}$: if $x=0$, Alice opens Charlie’s lab, and sets her outcome $a$ equal to his outcome $c$; if $x=1$, Alice unitarily reverses Charlie’s measurement process (i.e., applies unitary $\mathcal{U}^{\dagger}_{C}$ to lab $\mathbf{L}_{C}$) before opening Charlie’s lab, and performs some measurement of her own on the system, setting her outcome $a$ as the result of her measurement.
After Alice sets her outcome $a$, the system Charlie (if $x=0$) or Alice (if $x\neq 0$) just interacted with is sent into Debbie’s lab, where Debbie measures the system and gets the measurement result $d$. This measurement is characterised by the unitary $\mathcal{U}_{D}$ from outside of the lab. After Debbie’s measurement, the super-observer Bob acts based on his random input $y$: if $y=0$, Bob opens Debbie’s lab, setting his measurement outcome $b$ to be equal to Debbie’s outcome $d\in\{0,1\}$; whereas if $y=1$, Bob applies $\mathcal{U}^{\dagger}_{d}$ to unitarily reverse Debbie’s measurement of the system, before opening Debbie’s lab and performing a measurement of his own on the system, setting his outcome $b$ equal to the measurement result.
III.2 The Causal Friendliness No-Go Theorem
Building upon our description of Truly-Observed Events and Pseudo Events, and inspired by Refs. [17, 16], we first define:
Definition 1 (Operational Time Symmetry).
A prepare-measure experiment with preparation $\mathbb{P}$ and measurement $\mathbb{M}$ has operational time symmetry if there exists a time-reversed experiment, with preparation $\mathbb{P}^{{}^{\prime}}$ and measurement $\mathbb{M}^{{}^{\prime}}$, in which $\mathbb{P}^{{}^{\prime}}\equiv\mathbb{M}$ and $\mathbb{M}^{{}^{\prime}}\equiv\mathbb{P}$ such that,
$p_{\leftarrow}(f,e|v,u)=p_{\rightarrow}(e,f|u,v)\ \forall u,v,x,y.$ (6)
where $e$ and $f$ are outcomes corresponding to observed events when the measurements $u$ and $v$ respectively are performed.
Less formally, if an experiment has the property of Operational Time Symmetry, this means that if the roles of preparations and measurements in the experiment are swapped, reversing their temporal order, the observed probabilities should remain unchanged. The temporal ordering of events is represented in the causal influence diagram shown in Fig. 3 for better understanding.
A natural question is whether an operational theory should be expected to be Operationally Time Symmetric. The intuitive answer is No. This becomes clear if we recall the basic prepare-and–measure setup relevant to a laboratory experiment. Due to the thermodynamic arrow of time, it is natural to demand that future measurements should not affect past preparations, i.e.,
$p(e|u)=\sum_{f}p(e,f|u,v)=\sum_{f}p(e,f|u,v^{\prime})\quad\forall v,v^{\prime}.$ (7)
In contrast, it is not natural to expect the reverse condition, namely that preparations should not influence future measurements, i.e.,
$p(f|v)=\sum_{e}p(e,f|u,v)=\sum_{e}p(e,f|u^{\prime},v)\quad\forall u,u^{\prime},$ (8)
Figure 3: Causal influence diagram for the Causal Friendliness scenario. The rectangular entities are those that are in direct control of the observer, while the circles represents events that are not in direct control. The arrows represents the order of the events conveying the cause-effect relationships between the events. Red coloured shapes indicate the Pseudo Events while the blue one are Truly-Observed Events.
The impossibility of satisfying Eq. (8) is the main obstacle to having an Operationally Time Symmetric description of a theory. Nevertheless, without loss of generality, we may restrict our attention to those preparations for which Eq. (8) does hold. We refer to the corresponding set of correlations as no-signalling in time, denoted by $\mathcal{NS}_{T}$. It was shown in [17] that the sector of quantum theory satisfying $\mathcal{NS}_{T}$ is Operationally Time Symmetric. In what follows, while establishing our main results, we restrict to quantum preparations belonging to $\mathcal{NS}_{T}$ so that Operational Time Symmetry is guaranteed.
It is important to note that the definition of Operational Time Symmetry pertains exclusively to events that are both real and operationally accessible. In contrast, the Causal Friendliness scenario includes not only these real events but also additional elements referred to as Pseudo Events. Consequently, in such a scenario, only an analogue of Operational Time Symmetry can be meaningfully defined. Motivated by this distinction, we employ this analogue to establish the main results of the paper.
In Causal Friendliness scenario, the pair Charlie–Alice acts before Debbie–Bob in forward time, making the former the “preparers” and the latter the “measurers”. In the time-reversed picture, however, Debbie–Bob acts before Charlie–Alice, thus reversing the roles of “preparer” and “measurer”. At first glance, it may seem that as Charlie acts before Alice at a particular site of the whole scenario, this pair $(a,c)$ constitute a preparation and measurement in itself, where Charlie effectively prepares the system and Alice measures subsequently. Similar conclusions can be made for the other site where Debbie and Bob acts. However, this is not true. As already discussed, the events corresponding to $c$ and $d$ are Pseudo Events, as they are either eventually erased, or they end up corresponding to the outcomes $a$ and $b$ respectively. Therefore, we can only take the pair of events $(c,a)$ as a preparation, and the pair $(d,b)$ as a measurement, rather than splitting this up onto $a$ and $c$, and $b$ and $d$. This means the counterpart of the definition of Operational Time Symmetry applicable to this scenario does not involve reversing the entire temporal sequence $c\rightarrow a\rightarrow d\rightarrow b$ to $b\rightarrow d\rightarrow a\rightarrow c$. Instead, it reverses the temporal order of the grouped preparation and measurement events, namely $(c,a)\rightarrow(d,b)$, into $(d,b)\rightarrow(c,a)$. Considering this, a generalisation of Operational Time Symmetry can be proposed when this definition is extended to include the Pseudo Events. This forms the assumption of Axiological Time Symmetry, the final assumption needed for us to generate a full Causal Friendliness No-Go Theorem:
Theorem 1 (Causal Friendliness No-Go Theorem).
Consider the scenario we give in Section III.1, with Charlie and Alice acting before Debbie and Bob. Let $a,b,c,d\in\{\pm 1\}$ denote the outcomes of Alice, Bob, Charlie, and Debbie, respectively, and let $x,y\in\{0,1\}$ denote the measurement settings of Alice and Bob. Assume:
  1. 1.
    Absoluteness of Observed Events (AOE): All observed events at a particular point on the space-time region characterised by measurement outcomes $a,b,c,d$ have definite values, irrespective of any event that occurred at any other point in the space-time. This means the joint probability $p(c,a,d,b|x,y)$ exists and the observed probability $p(a,b|x,y)$ can be recovered from it via marginalisation as,
    $p(a,b|x,y)=\sum_{c,d}p(c,a,d,b|x,y)\ \ \forall x,y,a,b.$ (9)
  2. 2.
    Axiological Time Symmetry (ATS): If we invert the order of the scenario (i.e., have Debbie and Bob acting before Charlie and Alice), this inversion preserves the structure of these joint probabilities, i.e., $p_{\leftarrow}(d,b,c,a|y,x)=p_{\rightarrow}(c,a,d,b|x,y)\ \forall x,y,a,b,c,d$.
  3. 3.
    No Retrocausality (NRC): Any future measurement choice $x(y)$ cannot influence the past outcomes corresponding to the Pseudo Events $c(d)$, and any observed event $a$ in the past cannot be influenced by the future Pseudo Event $d$.
  4. 4.
    Screening via Pseudo Events (SPE): Every Truly-Observed Event is screened off from all the non-input variables in its causal past by appropriate pseudo-events.
    In the current scenario, for Bob’s observed event, this assumption implies:
    p(b|c,x,a,d,y)=p(b|c,x,d,y)b,c,x,d,yp(b|c,x,a,d,y)=p(b|c,x,d,y)\ \ $\forall$ b,c,x,d,y (10)
Given these Assumptions, the observed correlations satisfy the inequality:
$\Big|\,\langle A_{0}B_{0}\rangle+\langle A_{0}B_{1}\rangle+\langle A_{1}B_{0}\rangle-\langle A_{1}B_{1}\rangle\,\Big|\;\leq 2\,$, (11)
where $\langle A_{x}B_{y}\rangle=\sum_{a,b=\pm 1}ab\,p(a,b|x,y)$ are the joint expectation values for the measurements, conditioned on the inputs $(x,y)$.
The variables $c$ and $d$, representing the Pseudo Events, impose constraints on Truly-Observed Events strictly weaker than those imposed by hidden variables in conventional no-go theorems. Hidden variables are typically assumed to encode complete information about a physical system and (at least in principle) enable an event-by-event description of its behaviour. In contrast, the SPE assumption endows $c$ and $d$ with only limited power, allowing them to influence solely those Truly-Observed Events lying within their future light cones, and screening the future events off from any past events that are already influenced by, or that are already influencing, these Pseudo Events. For example, if the Pseudo Event $c$ exists and is absolute, then the Truly-Observed Event $a$ must always depend on $c$, regardless of the choice of $x$. In particular, for $x=0$ one has $a=c$, while for other values of $x$, even if $c$ were erased and the laboratory restored to its prior state, the effect of any past information on $a$ could still be argued to be screened through the pre-existing value of $c$. In this sense, so long as there is the assumption (implicit in, but weaker than, AOE), that the Pseudo Events $c$ and $d$ are Absolute, then SPE emerges as a very natural assumption. We will discuss this implicit assumption further in Section IV.3, where we attempt to subdivide AOE to see which parts of it are actually necessary to generate a contradiction with quantum mechanics. While the implications of removing AOE as a whole have been considered elsewhere [33], we believe we are the first to consider the effect of sub-assumptions of AOE.
The No Retrocausality assumption emerges by analogy to the idea of retrocausality in hidden variable models [16, 34]. Retrocausality is one of the possible interpretations of the violation of the Statistical Independence assumption in hidden variable models - i.e., a way of interpreting the idea that there may be a correlation between the hidden variables $\lambda$ in the model, and the measurement choices [11, 35]. Given these measurement choices are typically defined as being “free”, it is specifically the interpretation that the hidden variable $\lambda$ is in some way causally dependent on these measurement choices. This is as opposed to the interpretation of the violation of Statistical Independence as Superdeterminism, which denies this definition of the measurement choices as “free”, and instead says that they can be causally dependent on the hidden variables, or the interpretation of this violation through counterfactual restriction, where instead simply certain values of hidden variable and certain choices of measurement simply cannot exist, without there necessarily being a causal direction to that prohibition [13]. Note that, despite retrocausality implying a causal dependence of the hidden variables on the measurement choices, this does not necessarily imply any backwards-in-time signalling from the point the choices were made - indeed, most modern versions of retrocausality are all-at-once models, where, despite causal dependency, the model is treated as a “block universe” structure, where time is just a dimensional parameter, rather than implying any sort of ordered priority.
The retrocausality we appeal to with one of the parts of our NRC assumption takes a similar form to this retrocausality, except the prohibited relation is a causal dependence of the (potentially-observable) past Pseudo Events on the future measurement choices, rather than the dependence of any hidden variables on the measurement choices. This is as an instance of NRC (namely Pseudo Events are independent future measurement choices) where the Pseudo Events act equivalently to hidden variables in the conventional description of retrocausality. The other part of NRC comes from the idea that a Truly-Observed Past event $a$ cannot also depend on any of the non input variables $d$ that are potentially erased depending on a free choice.
From this point onward, we will collectively refer to the conjunction of the assumptions ATS, NRC, and SPE as Causal Agency. We now prove that the inequality given in Eq. (11) follows from the assumptions AOE and Causal Agency stated below.
III.3 Proof of the Causal Friendliness No-Go Theorem
Causal factorisation admitting NRC: In the forward time order, Charlie measures his quantum state first (obtaining $c$), then Alice chooses $x$ and so sets a value for $a$, then Debbie measures her quantum state (obtaining $d$), and finally Bob chooses $y$ and so sets a value for $b$. Thus, the most general causal factorisation for this scenario that admits NRC can be written as,
$\begin{split}p&(c,a,d,b|x,y)=\\ &p(c)\;p(a|x,c)\;p(d|c,x,a)\;p(b|c,x,a,d,y).\end{split}$ (12)
Consequence of ATS: From Bayes’ rule we can write,
$\begin{split}p(a,d|c,x)&=p(a|x,c)p(d|c,x,a)\\ &=p(d|x,c)p(a|c,x,d)\end{split}$ (13)
Inserting this into Eq. (12) for the forward-time scenario where Charlie-Alice measures before Debbie-Bob, we get,
$\begin{split}p&{}_{\rightarrow}(c,a,d,b|x,y)=\\ &p(c)p(d|x,c)p(a|c,x,d)p(b|c,x,a,d,y).\end{split}$ (14)
Similarly, for the time-reversed scenario, we have,
$\begin{split}p&{}_{\leftarrow}(d,b,c,a|y,x)=\\ &p(d)p(c|y,d)p(b|d,y,c)p(a|d,y,b,c,x).\end{split}$ (15)
This yields two key conditional independences, proved explicitly as Lemma 1 and Lemma 2.
Lemma 1.
If AOE, ATS and NRC hold, then Charlie and Debbie’s joint outcome statistics are independent of Alice and Bob’s choice of measurements $x$ and $y$, i.e.,
$p(c,d|x,y)=p(c,d).$
Proof.
From the assumption of ATS we have,
$p_{\leftarrow}(d,b,c,a|y,x)=p_{\rightarrow}(c,a,d,b|x,y)$ (16)
Summing over $a$, and $b$, Eq. (14) and (15) can be written as,
$\begin{split}p_{\rightarrow}(c,d|x,y)&=p(c)p(d|x,c)=p(c,d|x)\end{split}$ (17)
and
$p_{\leftarrow}(d,c|x,y)=p(d)p(c|y,d)=p(c,d|y)$ (18)
respectively.
Finally, combining Eqs. (17) and 18 with Eq. (16) gives,
$p_{\rightarrow}(c,d|x,y)=p(c,d|y)=p(c,d|x)=p(c,d)$ (19)
Lemma 2 ((Alice(Bob) cannot depend on Bob’s(Alice’s) choice of measurement beyond $c,d$).
If AOE, ATS and NRC hold, then Alice’s(Bob’s) outcome aa($b$) is effectively screened off by the Pseudo Events in its causal past; that is, it cannot depend on Bob’s (Alice’s) choice of measurement, and hence,
$p(a|c,d,x,y)$ $=p(a|c,x,d)$ (20)
$p(b\mid y,c,x,d)$ $=p(b\mid y,c,d).$ (21)
Proof.
By using Lemma 1 we can now write the joint probability as,
p(c,a,d,b|x,y)=p(c,d)p(a|c,x,d)p(b|c,x,a,d,y),p(c,a,d,b|x,y)=p(c,d)p(a|c,x,d)p(b|c,x,a,d,y), (22)
and then summing over $b$ gives,
$p(c,a,d|x,y)=p(c,d)p(a|c,x,d).$ (23)
Now dividing both sides by $p(c,d|x,y)$ (which by Lemma 1 is equal to $p(c,d)$) we get,
$p(a|c,d,x,y)$ $=\frac{p(c,a,d|x,y)}{p(c,d|x,y)}=p(a|c,x,d)$ (24)
However, due to NRC we can always write $p(a|c,x,d)=p(a|c,x)$.
Similarly, for the time-reversed scenario, using the ATS assumption (for notational simplicity $^{\prime}\rightarrow^{\prime}$ in the subscript is omitted), we can write,
p(c,a,d,b|x,y)=p(c,d)p(b|d,y,c)p(a|d,y,b,c,x),p(c,a,d,b|x,y)=p(c,d)p(b|d,y,c)p(a|d,y,b,c,x), (25)
Then summing over $a$ and dividing both side by $p(c,d|x,y)$ finally gives,
$p(b|c,d,x,y)=p(b|c,d,y)$ (26)
We have now established two important results in the form of Lemmas 1 and 2, which will be used to prove the main result. At this stage, we insert the SPE assumption (i.e., p(b|c,x,a,d,y)=p(b|c,x,d,y)b,c,x,d,yp(b|c,x,a,d,y)=p(b|c,x,d,y)\ \ $\forall$ b,c,x,d,y) into the causal factorisation (Eq. (12)), then inserting Lemmas 1 and 2, and NRC assumption to the modified expression yields
$p(c,a,d,b|x,y)$ $=p(c,d)\;p(a|x,c)\;p(b|y,c,d)\,.$ (27)
Marginalising over $c$ and $d$, and writing the probabilities $p(c)p(d|c)=p(c,d)$, gives the observed joint probability distribution as
$p(a,b|x,y)=\sum_{c,d}p(c,d)\;p(a|x,c)\;p(b|y,c,d).$ (28)
For each fixed pair $(c,d)$, we can define conditional expectation values
$\begin{split}A_{x}^{c}&:=\sum_{a=\pm 1}a\,p(a|x,c)\in[-1,1],\\ B_{y}^{cd}&:=\sum_{b=\pm 1}b\,p(b|y,c,d)\in[-1,1],\end{split}$ (29)
and so correlators
$\begin{split}\langle&A_{x}B_{y}\rangle=\sum_{a,b=\pm 1}ab\,p(a,b|x,y)\\ &=\sum_{c,d}p(c,d)\,\Big[\sum_{a}a\,p(a|x,c)\Big]\Big[\sum_{b}b\,p(b|y,c,d)\Big]\\ &=\sum_{c,d}p(c,d)\,A_{x}^{c}\,B_{y}^{cd}.\end{split}$ (30)
For each $(c,d)$, we define the functional,
$\begin{split}S^{cd}&:=A_{0}^{c}B_{0}^{cd}+A_{0}^{c}B_{1}^{cd}+A_{1}^{c}B_{0}^{cd}-A_{1}^{c}B_{1}^{cd}\\ &=A_{0}^{c}(B_{0}^{cd}+B_{1}^{cd})+A_{1}^{c}(B_{0}^{cd}-B_{1}^{cd}).\end{split}$ (31)
Since $B_{y}^{cd}\in[-1,1]$, and
$\forall u,v\in[-1,1],\,|u+v|+|u-v|\leq 2$ (32)
we can maximise over $|A_{0}^{c}|\leq 1$, $|A_{1}^{c}|\leq 1$ to get
$|S^{cd}|\leq|B_{0}^{cd}+B_{1}^{cd}|+|B_{0}^{cd}-B_{1}^{cd}|\leq 2.$ (33)
Therefore, the observed value of the function $S$ is
$\begin{split}S&=\sum_{c,d}p(c,d)\;S^{cd}\\ &=\sum_{c,d}p(c,d)\,\Big[A_{0}^{c}B_{0}^{cd}+A_{0}^{c}B_{1}^{cd}+A_{1}^{c}B_{0}^{cd}-A_{1}^{c}B_{1}^{cd}\Big]\\ &=\langle A_{0}B_{0}\rangle+\langle A_{0}B_{1}\rangle+\langle A_{1}B_{0}\rangle-\langle A_{1}B_{1}\rangle,\end{split}$ (34)
so by Jensen’s inequality [36],
$|S|\;\leq\;\sum_{c,d}p(c,d)\,|S^{cd}|\;\leq\;\sum_{c,d}p(c,d)\times\,2=2.$ (35)
Quantum violation: To achieve the maximal violation of the inequality in Eq. 35, Charlie and Alice must perform measurements of complementary observables, as should Debbie and Bob. Under such choices, the quantum value reaches
$S_{\mathrm{QM}}=2\sqrt{2}>2.$
For example, one optimal choice of measurement settings is:
$\text{Charlie: }\sigma_{z},\quad$ $\text{Alice: }\sigma_{x},$
$\text{Debbie: }\frac{\sigma_{z}+\sigma_{x}}{\sqrt{2}},\quad$ $\text{Bob: }\frac{\sigma_{z}-\sigma_{x}}{\sqrt{2}}.$
Therefore, no theory satisfying AOE and Causal Agency can reproduce all quantum Causal Friendliness correlations. ∎
IV Role of Pseudo Events in the Causal Friendliness Paradox
From the analysis in the previous section, a natural question arises: does the violation of Eq. (35) imply that the observed events are not absolute (even the Pseudo Events), but instead subjective to the observers? To address this, we first examine to what extent the AOE assumption is truly necessary, in addition to ATS, NRC, and SPE, for establishing the Causal Friendliness no-go theorem. In this context, we attempt to supersede the AOE assumption by drawing an analogy with the replacement of the “wavefunctions are $\psi$-ontic” assumption used in Price’s no-go result [16], by the $\lambda$-mediation assumption (discussed later in this section), which Leifer and Pusey used to establish a no-go theorem [17] about time symmetry and retrocausal ontology. Price’s no-go theorem [16] suggests that any time-symmetric ontology of quantum theory must entail retrocausality. His argument proceeds by showing that the following three assumptions, when taken together, contradict quantum theory:
  • No Ontic Retrocausality: Future measurement choices cannot influence the distribution of past ontic outcomes. In the forward-time ordering this gives the most general causal factorisation
    $p(\lambda,a,b|x,y)\;=\;p(\lambda)\,p(a|x,\lambda)\,p(b|a,\lambda,y).$ (36)
    Note this is defined analogously to condition NRC given earlier in this paper, but applied to hidden variable $\lambda$ rather than Pseudo Events $c$ and $d$;
  • Ontological Time-Symmetry: The ontological probabilities for the forward ordering equal those of the appropriate time-reversed ordering, i.e.
    $p_{\rightarrow}(\lambda,a,b\mid x,y)\;=\;p_{\leftarrow}(b,a,\lambda\mid y,x).$ (37)
    This is the the ontological counterpart of the ATS assumption, that is applied to probabilities associated with $\lambda$ instead of the Pseudo Events;
  • Reproducibility with a $\psi$-ontic ontological model: Compatibility of quantum statistics with the assumption that “No two distinct quantum states have overlapping ontic state distributions”, in the sense of Harrigan and Spekkens’s definition of $\psi$-ontic [37], formulated within Spekkens’s ontological models framework [38].
Note however that ontological models are only a small subset of possible hidden variable models. Further, Ref. [39, 40] recently showed that ontological models have trouble reproducing key aspects of the predictions of quantum mechanics, even before we consider the subset which are $\psi$-ontic. Therefore, reproducibility with a $\psi$-ontic model is not a natural assumption to use when proving the fact that time-symmetric interpretation of quantum theory would be retrocausal. The status of the state, whether it is ontic or epistemic (or potentially both [41]), should not come into the proof. In order to address this problem, and to prove Price’s result without attributing any ontological status to the state, Leifer and Pusey [17] showed that combining Price’s assumptions of No Retrocausality and Time-symmetry with another assumption, $\lambda$-mediation, leads to a contradiction with quantum theory [17].
IV.1 Replacing $\psi$-ontic reproducibility with $\lambda$-mediation
In the ontological models framework, where the quantum state (wavefunction) is represented by a probability distribution over a space of underlying ’ontic’ states (see [42] for details), models are classified as either $\psi$-ontic or $\psi$-epistemic. In $\psi$-ontic models, distinct quantum states correspond to non-overlapping distributions over this ontic state space. $\psi$-epistemic models instead allow quantum states to overlap on this space. The extra assumptions imposed on the possible ontological counterpart of the quantum state often restrict the statistics that such state can generate.
Price’s result [16] on the time-symmetric ontology of quantum theory lacks generality, since his argument relies explicitly on assuming reality is $\psi$-ontic. While $\psi$-epistemic models are themself subject to significant criticisms [43, 44, 45, 46, 47, 48, 49, 50], even together these two categories leave out all models of the wavefunction which do not follow the ontological models framework. This suggests that, to properly investigate time-symmetric ontologies, we should avoid limiting ourselves to one or the other categories of ontological model.
In this direction, Leifer and Pusey [17] proposed a weaker requirement, known as $\lambda$-mediation, according to which all correlations between preparations and measurement outcomes are mediated via the underlying ontic state $\lambda$ in an Ontological Model of the scenario. Formally, this means that the probability of an outcome $b$ for a arbitrary preparation $P$ and a measurement $M$ can be written as
$p(b|P,M)=\sum_{\lambda}p(\lambda|P)\,p(b|M,\lambda),$ (38)
without requiring that the ontic distributions $p(\lambda|P)$ corresponding to different quantum states be either disjoint or overlapping. Thus, $\lambda$-mediation relaxes Price’s $\psi$-ontic reproducibility assumption by removing the requirement of attributing a particular status to quantum state, while still preserving the idea that $\lambda$ screens off the correlations between preparation and measurement.
We here ask whether we can relax the AOE assumption to a weaker assumption in the same way as Leifer and Pusey replace Price’s $\psi$-ontic assumption with $\lambda$-mediation [17]. Given the AOE assumption seems similar to the $\psi$-ontic assumption (albeit imposed on the operational and observational level instead of the ontological/hidden-variable level), we argue that the answer is “yes”. To show this is the case, we first examine formally the relation between the AOE assumption, and the $\psi$-ontic criterion for ontological models. We then show that the AOE assumption can be relaxed, to an operational equivalent of $\lambda$-mediation, which is weaker than AOE, while still allowing us to form Local Friendliness-style inequalities which quantum theory violates.
IV.2 Can we weaken the AOE assumption?
The logical structure underlying our weakening of the AOE assumption closely follows Leifer and Pusey’s weakening of Price’s $\psi$-ontic assumption to $\lambda$-mediation [17] - both replacing a strong absoluteness assumption ($\psi$-ontic reproducibility or AOE) with a weaker one (ontic or operational mediation). The key similarity between the arguments in [17] and ours lies in the fact that the independence of the mediator variables from the choice of measurement settings are not assumed but rather derived from other underlying assumptions.
Theorem 2 (Operational Pseudo Event Mediation (OPEM)).
Let the Causal-Friendliness scenario be as in Sec. III. We replace AOE (existence of the full joint $p(c,a,d,b|x,y)$) with the following weaker, purely operational assumptions:
  1. 1.
    Existence of marginals (EOM): For all settings $(x,y)$ there exist well-defined operational marginals $p(c,d|x,y)$ and conditional response functions $p(a|x,c)$ and $p(b|y,c,d)$ describing how the outcomes of Truly-Observed Events $a$ and $b$ depend on Pseudo Events and choice of measurement settings. (Note we make no claim about existence of the full four-variable joint probability for each run.)
  2. 2.
    Operational mediation (OM): Pseudo Events $c,d$ acts as the mediator between future and past events such that once $c,d$ are specified, $a$ provides no further information on $b$ beyond that already provided by $c,d$ and $y$, i.e.,
    $p(b|a,c,d,x,y)=p(b|c,d,y),$ (39)
    for all arguments.
Combining these with the ATS and NRC assumptions (as in Sec. III), then the operational probabilities admit the factorisation
$p(a,b|x,y)\;=\;\sum_{c,d}p(c,d)\,p(a|x,c)\,p(b\mid y,c,d),$ (40)
and hence the Causal-Friendliness bound $|S|\leq 2$ follows.
Proof.
In order to prove the the factorisation of operational probability $p(a,b|x,y)$ as given in Eq. (40), we recall the operational ingredients posited by OPEM - specifically,
$p(b|a,c,d,x,y)=p(b|c,d,y)\qquad\forall a,b,c,d,x,y.$ (41)
Crucially, we do not assume the prior existence of a Kolmogorovian joint probability measure $p(a,b,c,d\mid x,y)$ defined by all the four events. Instead construct the probability from the operational marginals that are empirically defined. To this end we define such a operationally constructed distribution for operational joint probability-like object as
$q(a,b,c,d|x,y)\;:=\;p(c,d|x,y)\;p(a\mid x,c)\;p(b|y,c,d).$ (42)
The right-hand side of Eq. (42) uses only the operational objects that OPEM assumes exist. Therefore, $q$ should be regarded as a constructed function of these empirical distributions, rather than as a postulate implying the prior existence of a full joint distribution. However, we expect that for every fixed $(x,y)$, $q(\cdot,\cdot,\cdot,\cdot|x,y)$ is a non-negative function which sums up to unity as each factor constitutes a conditional probability:
$\sum_{a,b,c,d}q(a,b,c,d|x,y)$
$=\sum_{c,d}p(c,d|x,y)\Big(\sum_{a}p(a\mid x,c)\Big)\Big(\sum_{b}p(b|y,c,d)\Big)$
$=\sum_{c,d}p(c,d|x,y)=1.$ (43)
Hence $q(\cdot,\cdot,\cdot,\cdot\mid x,y)$ is a valid probability distribution. Defining $q$ in Eq. (42) is not the same as assuming the existence of the previously given Kolmogorovian joint probability. Rather, $q$ is built from marginals and conditionals which are assumed to be empirically defined entities; this is a strictly weaker commitment than postulating a global joint probability a priori.
Now the statistics corresponding to the Truly-Observed Events $a$ and $b$ can be recovered by marginalising $q$ as,
$\sum_{c,d}q(a,b,c,d|x,y)$
$=\sum_{c,d}p(c,d|x,y)\,p(a|x,c)\,p(b|y,c,d)$
$:=\widetilde{p}(a,b|x,y).$ (44)
By construction $\widetilde{p}(a,b|x,y)$ is the operational prediction obtained by combining the assumed marginals in OPEM. To prove that $\widetilde{p}(a,b|x,y)$ is same as $p(a,b|x,y)$ given in Eq. (40) we have to prove the independence of $c,d$ from $x,y$. However, we cannot establish this result in the same manner as Lemma 1, since the proof of Lemma 1 relies on the assumption of AOE, which we do not use here. Using instead the operational mediation assumption of OPEM, we write the constructed distribution as in Eq. (42).
Further, ATS implies that the forward and time-reversed operational constructions coincide under the exchange $(c,a,x)\!\leftrightarrow\!(d,b,y)$, meaning we can write,
q(a,b,c,d|x,y)=p(c,d|y,x)p(b|y,d)p(a|x,c,d),q(a,b,c,d|x,y)=p(c,d|y,x)\,p(b|y,d)\,p(a|x,c,d), (45)
where the quantities on the right-hand side are again operational marginals defined for the time-reversed experiment. Note that, every object that appears in the algebra above (namely, the factors in $q$ and the equalities given by ATS) is either an operational marginal (e.g. $(p(c,d|x,y)$) or an operational conditional (e.g. $p(a|x,c)$, $p(b|y,c,d)$), all assumed to exist by OPEM. We never invoked any property that would require these to be derived from the same pre-existing global Kolmogorovian probability distribution; rather we constructed $q$ from the operational probabilities. Equating Eqs. (42) and (45) gives
$\begin{split}p(c,d|x,y)\,&p(a|x,c)\,p(b|y,c,d)\\ &=p(c,d|y,x)\,p(b|y,d)\,p(a|x,c,d).\end{split}$ (46)
Summing over $a$ and $b$ we get from Eq. (46) $p(c,d|x,y)=p(c,d|y,x)$. Finally, applying NRC in the forward and time reversed scenario, we get $p(c,d|x,y)=p(c,d|x)$ and $p(c,d|y,x)=p(c,d|y)$. This simply implies that they both are equal to $p(c,d)$, completing the proof. ∎
It is important to note that the assumption OM is stronger than the SPE assumption used in the proof of the Causal-Friendliness no-go theorem. In other words, SPE can be viewed as a specific instance of Operational mediation, where the dependence of the event $b$ on the past input choice $x$ is not relaxed. In the proof of Causal-Friendliness no-go theorem, however, the independence of $b$ from $x$ is derived from the joint assumptions of AOE, ATS, and NRC. In contrast, in the present analysis, we are not permitted to assume AOE at the outset, and consequently, the independence of $b$ from $x$ is incorporated directly into the relatively stronger assumption of Operational mediation itself.
Another important point to note here is that, although EOM by itself does not require the existing marginals to originate from a single global joint probability distribution, imposing the requirement that those existing marginals given by EOM originate from a single global joint probability distribution leads to AOE. This additional requirement can be summarised as follows:
Definition 2 (Existence of Joint Probability Distributions (EJPD)).
A theory is said to satisfy the Existence of Joint Probability Distributions if all marginals corresponding to the observed events can be recovered from a single global joint probability distribution $p(a,b,c,d|x,y)$.
Then by considering the above definition we have,
$\textbf{EOM}+\textbf{EJPD}\implies\textbf{AOE}$ (47)
IV.3 Impossibility of discarding the Absoluteness of Pseudo Events
Above, we showed that we can split the AOE assumption into EOM and EJPD, and still obtain a contradiction with quantum mechanics if we replace AOE with EOM, so long as we strengthen SPE into OM. We now argue that we can split EOM up further, into Absoluteness of Truly-Observed Events ATOE, and Absoluteness of Pseudo Events APE defined respectively as,
Definition 3 (Absoluteness of Truly-Observed Events (ATOE)).
A theory satisfies Absoluteness of Truly-Observed Events if all observed events at a particular point on the space-time region characterised by measurement outcomes $a,b$ have definite values regardless irrespective of any event that occurred at any other point in the space time. This means that the joint probability $p(a,b|x,y)$ exists.
Definition 4 (Absoluteness of Pseudo Events (APE)).
A theory satisfies Absoluteness of Pseudo Events if all observed events at a particular point on the space-time region characterised by measurement outcomes $c,d$ have definite values regardless irrespective of any event that occurred at any other point in the space time. This means that the joint probability $p(c,d|x,y)$ exists.
We can summarise the connections between the assumptions discussed in Sec. IV.3 and Sec. IV.2 as follows:
$\begin{split}\textbf{EOM}&+\textbf{OM}\implies\textbf{OPEM}\\ \textbf{EOM}+\textbf{NRC}&\implies\textbf{ATOE}+\textbf{APE}\end{split}$ (48)
The second implication in Eq. (48) is due to the fact that the assumption EOM already assumes that $p(c,d|x,y)$ exists ( i.e.,APE). Moreover, the second part of EOM is that $p(a|x,c)$ and $p(b|y,c,d)$ exists. Then together this implies that there exists a probability $p(a,b|x,y)$ as
$p(a,b|x,y)$ $=\sum_{c,d}p(c,d|x,y)p(a|x,c)p(b|y,c,d)$ (49)
$=\sum_{c,d|x,y}p(c,d)p(a|x,c,d)p(b|y,c,d)$
which is what ATOE is. The second line of the above equation is due to NRC.
Having shown in Section IV.2 that we can remove the EJPD assumption while still being able to generate an inequality which quantum mechanics can violate, we now wish to consider whether we can further remove assumption, ideally to see if ATOE alone is enough to produce an inequality which Quantum Mechanics violates. This question is motivated by prima facie APE being violated by standard quantum mechanics, given quantum mechanics describes Charlie’s (Debbie’s) lab as being in a superposition from Alice’s (Bob’s) perspective until they open the lab, and given Alice (Bob) can unitarily undo Charlie’s (Debbie’s) measurement - which requires the measurement itself to have been unitary according to quantum mechanics in the first place.
To answer this question, let us consider the case without APE, where we have the following factorised form of observed probabilities,
$p(a,b|x,y)=\sum_{c,d}p(c,d|x,y)\,p(a|x,c)\,p(b|y,c,d).$ (50)
which gives,
$\begin{split}&S=\\ &\sum_{c,d}\Big[p(c,d|0,0)A_{0}^{(c)}B_{0}^{(c,d)}+p(c,d|0,1)\,A_{0}^{(c)}B_{1}^{(c,d)}\\ &+p(c,d|1,0)A_{1}^{(c)}B_{0}^{(c,d)}-p(c,d|1,1)\,A_{1}^{(c)}B_{1}^{(c,d)}\Big],\end{split}$ (51)
We can no longer factor this as a single convex average $\sum_{c,d}p(c,d)\,S_{c,d}$ as done while establishing inequality Eq. (35). The pointwise bound $|S^{c,d}|\leq 2$ for fixed $c,d$ does not, by itself, constrain mixtures with different context-dependent weights. As a result, the overall value of $S$ derived without assuming APE may even exceed the quantum bound, undermining the no-go theorem.
To prove this, consider the specific case where $c,d\in\{0,1\}$ and $A_{x}^{(c)},B_{y}^{(c,d)}\in\{\pm 1\}$ the deterministic responses and set
[…]
Now if we choose responses such that $A_{0}^{(0)}=A_{1}^{(1)}=B_{0}^{(0,0)}=B_{1}^{(0,1)}=B_{0}^{(1,0)}=+1$ and $B_{1}^{(1,1)}=-1$, then,
$\langle A_{0}B_{0}\rangle=\langle A_{0}B_{1}\rangle=\langle A_{1}B_{0}\rangle=+1,\langle A_{1}B_{1}\rangle=-1.$ (53)
which implies S=1+1+1(1)=4S=1+1+1-(-1)=4, violating not only the bound provided in Eq. (35) but also the quantum bound $S_{Q}=2\sqrt{2}$, reaching the Boxworld bound $S_{B}=4$ [18]. This shows that without mediator independence (derived by combining ATS+NRC (Lemma 1) with AOE), it is not possible to establish the no-go result.
V Discussions and outlook
V.1 Comparing to Related Results
Recent works have connected Local Friendliness–type no-go theorems to both forms of contextuality, namely preparation contextuality [51] and Kochen–Specker contextuality [52]. These works demonstrate that different manifestations of quantum correlations can be used to establish the incompatibility of quantum theory with observer-independent events. This relationship can be understood in terms of a hierarchy of correlations that each no-go theorem employs.
In a prepare-and-measure scenario the connection between preparation contextuality and AOE is formulated through what they refer to as the Operational Friendliness no-go theorem. Crucially, Leifer and Pusey proved that, in a timelike prepare–and–measure scenario, imposing Time Symmetry together with No Retrocausality and $\lambda$-mediation yields the mediator/epistemic independence condition (e.g., $p(\lambda|x)=p(\lambda)$ for the no-signalling preparations) usually assumed in preparation-noncontextuality arguments [17]. Consequently these assumptions together imply a timelike Bell-CHSH constraint which quantum theory violates. In contrast, the Operational Friendliness no-go theorem derives this contradiction with quantum theory by combining AOE with Operational Agency- a stronger version of operational equivalence used in the discussion of Spekkens’ contextuality[38, 51, 52]. By using similar operational correlations involving time-ordered measurements, we employ AOE to establish our Causal Friendliness no-go result through a different yet equally powerful route. We therefore view our Causal Friendliness no-go theorem as complementary to the Operational Friendliness no-go theorem. However, the two are not hierarchically ordered because Time Symmetry and preparation-noncontextuality are logically different principles and can be independent in other scenarios.
Note that the Causal Friendliness constraints are less strict than the generalised-noncontextuality constraints for the sequential measurement scenario. However, for the bipartite scenario we discuss here, where each party can choose between two dichotomic measurements (the $2222$-Bell scenario), noncontextuality constraints provide the same Bell-CHSH-like inequality. To prove that our constraint is strictly weaker than those noncontextual constraints, one would need to consider scenarios beyond the $2222$-Bell scenario, where one can violate a Causal Friendliness inequality but not an inequality derived constraints similar to those used in proofs of Kochen-Specker contextuality [52] or preparation contextuality [51]. We leave the derivation of such Causal Friendliness inequalities beyond the $2222$-Bell scenario for future work.
V.2 Implications of the findings for interpretational issues
It is worth examining how the Absoluteness assumptions we discuss above relate to two prominent realist interpretations of quantum theory: the de Broglie–Bohm (dBB) theory [53, 54, 55], and Everett’s Many–Worlds interpretation (MWI) [56].
In the dBB framework, every system is associated with a set of ontic variables (the system’s exact position) that evolves under a guiding equation. These variables assign definite outcomes to all measurement events, so both the Truly-Observed Events and the Pseudo Events possess ontic definiteness in every run. In the terminology of this work, dBB explicitly maintains ATOE and APE. However, as the quantum potential is explicitly nonlocal (allowing measurement results to depend on factors outside the past light-cone of the measurement), in the Time Symmetric sector, dBB circumvents our no-go results by abandoning the condition of SPE.
On the other hand, the Many-Worlds Interpretation (MWI) fundamentally rejects the very notion of single, absolute outcomes. In the MWI, a measurement does not yield a unique, classical result. Rather, it induces a branching of the global wavefunction into dynamically independent components, each realising a distinct record of the possible outcomes. Consequently, the MWI explicitly denies the AOE assumption on which our no-go theorem is based. From the perspective of an MWI adherent, the apparent incompatibility between quantum predictions and the Causal-Friendliness inequality simply reflects the fact that the observed events in any given branch represent only one relative component of a larger, branch-dependent reality. In this view, the violation of AOE at the operational level arises naturally from the multiplicity of branching records in the universal wavefunction, thereby making the Causal-Friendliness no-go result completely compatible with the logic used in MWI.
V.3 Summary and future directions
In this paper, building on recent results, we introduced the Causal Friendliness Paradox, a time-ordered analogue of Bong et al ’s Local Friendliness no-go theorem. To generate this no-go theorem, we replace the usual locality assumption with Axiological Time Symmetry (ATS), showing that, when combined with Absoluteness of Observed Events (AOE), No Retrocausality (NRC), and Screening via Pseudo Events (SPE), we obtain a causal inequality. We then showed that quantum mechanics violates this inequality, demonstrating quantum statistics must be incompatible with at least one of these assumptions. To probe which assumption might be incompatible, we then examine whether AOE in its entirety is essential for this no-go result. We propose a weaker, operational form of AOE that still leads to the inequality that quantum mechanics violates. This result shows that even under relaxed assumptions, quantum theory resists reconciliation with classical notions of absolute events, reinforcing the foundational significance of Wigner’s Friend-type paradoxes in timelike scenarios. However, we finally showed that trying to remove the absoluteness of the individual Pseudo Events $c$ and $d$ undermines our ability to arrive at a contradiction with quantum mechanics, allowing us to not only surpass the maximum quantum value, but reach the maximum possible algebraic value for the correlators for the scenario. This illustrates the crucial role played by the Absoluteness of Pseudo Events in establishing the no-go result.
Our findings opens several directions for the future work, such as extending the weakening the Absoluteness of Observed Events to derive other no-go theorems based on different set of assumptions (e.g. macrorealism [57]), as well as exploring experimentally accessible scenarios that could test genuine Causal-Friendliness violations. A further direction is to relate the present framework to studies of the quantum-to-classical transition. Since the framework of weakening of absoluteness of events specifies when events acquire effective classicality, it may offer an operational perspective on when and how definite outcomes emerge through decoherence [58, 59] and coarse-grained measurements [60].
Acknowledgements - SM thanks Alok Kumar Pan for many insightful discussions on Wigner’s Friend paradoxes over the past years. JRH acknowledges support from a Royal Society Research Grant (RG/R1/251590), and from their EPSRC Quantum Technologies Career Acceleration Fellowship (UKRI1217).